OpenStudy (anonymous):

A beaker contains 40 g of water at 20 degree C. Now 50 g of ice is put into the beaker. The resulting temperature will be

OpenStudy (anonymous):

this problem is done assuming that all ice melts to mix with water first u need to know specific heat of water=4200J/kgK The specific heat is the amount of heat per unit mass required to raise the temperature by one degree Celsius. heat gained by water=heat lost by ice heat gained by water=ms(T2=T1) s is specific heat heat lost by ice=Q=mL+ms(T2-T2') T2'=o degrees(starting temp of moten water) where l is latent heat of ice(heat req to change it from solid to liquid sate) we have heat req to change the stae of ice+heat req to raise temp of melted water(as final temp is not 0degrees celcius) L=335J/g (latent heat of water) equate both to find T2